OpenStudy (anonymous):

i need help with factoring 81x^6+24x^3y^3 (the answer is 3x^3(3x+2y)(9x^2-6xy+4y^2) )

OpenStudy (anonymous):

please?

OpenStudy (anonymous):

first find the GCF of the two terms this is 3x^3 so we get 3x^3(27x^3 + 8y^3) now we use the identity x^3 + y^3 = ( x + y)( x^2 - xy + y^2) replace x by cube root of 27 x^3 and y by cube root of 8 y^3 which is 3x and 2y respectively and you'll obtain the answer you want

OpenStudy (anonymous):

are you ok with that?

OpenStudy (anonymous):

how would i find the cube root of 27x^3 and 8y^3

OpenStudy (anonymous):

i've given you them 27 x^3 : cube rt 27 = 3 and of x^3 = x :- 3x

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

when you say, "now we use the identity," what does that mean?

OpenStudy (anonymous):

nevermind. i got it :)

OpenStudy (anonymous):

well 27x^3 + 8y^3 is the sum of two cubes as is x^3 + y^3

OpenStudy (anonymous):

oh ok - good

OpenStudy (anonymous):

plz click on good answer if you are happy with these answers